Leslie is 5 times Rebecca's age. In 2 years, Leslie will be 3 times Rebecca's age. How old is Rebecca now?
Answer by rapture(86)

Leslie's current age = 5x

Rebecca's current age = x

Leslie's age is 2 years = (5x + 2)

Rebecca's age in 2 years = x + 2

Here is your equation:

5x + 2 = 3(x + 2)

Solve for x to find your answer.